OpenStudy (anonymous):

Please HELP ME I WILL GIVE A MEDAL!!!! 4x-y=17 4x+5y=-37 what is the order pair? consistent or not? dependent or not?

OpenStudy (anonymous):

The two lines are not parallel. Therefore there is only one solution and the system is consistent and independent

OpenStudy (anonymous):

WHAT ABOUT THE ORDER PAIRS? PLEASE

OpenStudy (anonymous):

There is only one solution. To find it you can simply subtract one of the equations from the other to eliminate a variable and solve for the other one.

OpenStudy (anonymous):

17,-37

OpenStudy (anonymous):

To find the solution, Subtract equation (2) from equation (1), you get: 4x-y=17 - 4x+5y=-37 --------------- -6y=54 ---> y=-9 Substitute this value of y in equation (1), you get: \[4x+9=17 \implies 4x=8 \implies x=2\] That's the solution is (2,-9)